OpenStudy (anonymous):

give me 3 points of this equation -3y= -x -6

OpenStudy (anonymous):

Two points can be found at the x and y intersections. To make things easiest the first step would be to simplify the equation to point inter-slope form by dividing both sides by -3. y=-(1/3)x+2 Then you would plug in 0 for x and solve, and then plug in 0 for y and solve. Don't forget the points are listed in (x,y) form. Solving those would give you two points and then a third point would be x=3 and solving for y. Here is the three examples just plug in the solutions for the letters: (0,y) (x,0) (3,y)

OpenStudy (anonymous):

i still dont get it

OpenStudy (anonymous):

Your given equation is: -3y= -x -6 my first step is to divide both sides by -3: \[-3y/-3 = (-x-6)/-3\] then if you split it into two parts on the left side to simplify more \[y=(-x/-3)+(-6/-3)\] the negatives cancel out for the -6/-3 \[y=-(x/3)+2\] now you can plugin numbers to find the points: first we will do x=0 to find a value for y. y=-(0/3)+2 the 0/3=0 leaving only a 2 so y=2 that makes the first point in x,y form (0,2) let me know if you still need help

OpenStudy (anonymous):

can you help with the second an third one?

terenzreignz (terenzreignz):

While that may have made it look a little harder than it should, there is no such concept as THE second and third one, they all depend on you, actually. Let's run that by you again: -3y= -x -6 To get ANY point on this line, pick any three DIFFERENT values for x. Say x = 0 Replace x with 0. -3y= -0 -6 And just solve for y. -3y = -6 Divide both sides by -3, and you should get y = -6/-3 y = 2 Hence, when x = 0, y = 2. Therefore, the point (0,2) is on the line.

OpenStudy (anonymous):

can you do that with a different x like 2?

terenzreignz (terenzreignz):

Any value for x, in fact. So, pick another value for x, one that isn't 0.

OpenStudy (anonymous):

2

terenzreignz (terenzreignz):

Fine... -3y= -x -6 Replace x with 2. -3y = -2 - 6 and solve for y.

OpenStudy (anonymous):

thats what i dont get is solving for y

terenzreignz (terenzreignz):

You can't solve linear equations in one variable? D:

OpenStudy (anonymous):

no i do i just dont know what to do

terenzreignz (terenzreignz):

Well, then solve -3y = -2 - 6 as though it were a linear equation, then.

terenzreignz (terenzreignz):

Heck, it is....

OpenStudy (anonymous):

what??

terenzreignz (terenzreignz):

You said you can solve linear equations in one variable... then solve -3y = - 2 - 6 Because that ^ IS a linear equation in one variable.

OpenStudy (anonymous):

what do i do first?

terenzreignz (terenzreignz):

Let me run the first one by you again... with detailed steps...

terenzreignz (terenzreignz):

\[\large -3y = -x -6\]Now suppose x = 0. Replace x with 0. \[\large -3y = -\color{red}0 - 6\] Now simplify -0-6, they are just numbers which may simply be added or subtracted... \[\large -3y = \color{green}{-6}\] Now divide both sides by -3. \[\large \frac{-3y}{\color{blue}{-3}}=\frac{-6}{\color{blue}{-3}}\] This cancels:\[\large \frac{\cancel{-3}y}{\color{blue}{\cancel{-3}}}=\frac{-6}{\color{blue}{-3}}\] Leaving you with \[\color{red}{\huge y = 2}\] And so, the point \[\huge (0,2)\] is one point on your line. Now please do the same, starting from x = 2.